OpenStudy (anonymous):

What is 4x + 2y = -20 written in slope-intercept form?

OpenStudy (anonymous):

Y = -2X - 10. Basically; you're re-writing the equation so it's in slope-intercept form (Y=MX+B) so that you can get your answer.

OpenStudy (anonymous):

do you think you can answer three more questions? Ill become fan!

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Given the points (2, 3) and (6, 8), write the point-slope form of the equation.

OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line which passes through the points (1, 6) and (2, 9).

OpenStudy (anonymous):

Write the equation y = 2 - 4x in standard form.

OpenStudy (anonymous):

#1: Point slope form is Y-Y1=M(X-X1). So we can say that the answer is: Y - 3 = 5/3 (X-2).

OpenStudy (anonymous):

#2: Slope Intercept for is Y=MX+B so we can use the following: Y = 3X+3